OpenStudy (anonymous):

Find the quotient using the rules for exponents: 8^6 divided by 8^2 8x8 8^4 8^3 8^8

OpenStudy (anonymous):

a^x * a^y = a^(x+y) a^x / a^y = a^(x-y) 8^6 / 8^2 = 8^(6-2) = 8^4